The farmer produces 102 bushels of wheat at a total cost of $2 per bushel. He sells all of the wheat to Firm F for $4 per bushel.

Firm F produces 59 pounds of flour from the wheat at a total cost of $6 per pound (including the amount paid to the farmer for the wheat). Firm F sells 43 pounds of flour to consumers for $12 per pound.

There are no other firms in this simple economy.

In total, how much profit do the farmer and Firm F earn?

Enter a whole number with no dollar sign.
Answer #1

Answer: $366

The farmer produces 102 bushels of wheat at a cost of $2 per bushel.

The total cost of 102 bushels = $2 * 102 = $204

He sells all of the wheat to Firm F for $4 per bushel.

The total revenue the farmer earns = $4 * 102 = $408

The total profit the farmer earns = Total revenue - Total cost

Or, the total profit the farmer earns = $408 - $204 = $204

So, the farmer earns the total profit of $204.

The firm F produces 59 pounds of flour from the wheat at a total cost of $6 per pound.

The total cost of production of 59 pounds of wheat = $6 * 59 = $354

The firm F sells 43 pounds of flour to consumers for $12 per pound.

The total revenue the firm F earns from selling wheat = $12 * 43 = $516

The total profit the farm F earns = Total revenue - Total cost

Or, the total profit the farm F earns = $516 - $354 = $162

The total profit the farmer and farm F earn = $204 + $ 162 = $366

______________________________________________________________________